# FlexiBit: Fully Flexible Precision Bit-parallel Accelerator Architecture for Arbitrary Mixed Precision AI

### v1

@@ 5.3.1. Latency Comparison
Figure 15, presents the latency evaluation results of FlexiBit and baselines. We observe no significant improvements in FP16-based models. However, as we deploy non-power-of-two prefcisions, FlexiBit significantly outperforms all the baselines, providing 59% and 31% less latency compared to Tensor Core and Bit-Fusion, on average. The improvements of FlexiBit mainly originate from the capability to support arbitrary precision with high hardware underutilization. Baselines need to up-cast non-power-of-two precision data to the nearest power-of-two precision data, which leads to significant redundant bit operations. Another source of improvement is our $BitPacking$ technique which fully utilizes the SRAM capacity, which helps in reducing DRAM traffic.
Figure 16. Energy comparison on 11 Act./Wgt. precision pairs normalized against TensorCore-like accelerator. The sequence length is 2048.
5.3.2. Energy Comparison
@@
Figure 16 presents the energy comparision results. We first observe that FlexiBit with weight-stationary dataflow provides considerably low energy, 66% and 33% lower than the Tensor Core and Bit-Fusion baselines, on average. The main source of the energy saving is reduced on-chip traffic by utilizing packed memory layout. As a result, the energy saving is significant when non-power-of-two precision data is involved. However, we observe that FlexiBit with weight-stationary involves small overhead for 16-bit arithmetic in some cases, which accounts for the static costs of the flexibility.
We also observe output stationary dataflow results in significant energy consumption increase on FlexiBit. This is due to the lack of input and weight reuse of the output-stationary dataflow we adopted, while the output reuse is restricted by the bitwidth of PE datapath and small output buffer size. The results indicates that dataflow-architecture co-design can significantly impact the energy efficiency of flexible bit-precision accelerator architecture, which implies further optimization opportunity in the design space.
@@
Figure 17. Performance per area comparison on 11 Act./Wgt. precision pairs. The sequence length is 2048.
@@
Figure 18. Comparison of bit-serial (Cambricon-P) and -parallel (FlexiBit) flexible precision architecture using EDP normalized against Tensor Core-like.
@@
Figure 19. Area breakdown of FlexiBit on Mobile-A scale and Reg_Width = 16(see table 3 for hardware configurations).
@@ 5.3.3. Performance Per Area
To understand the relative benefits over costs, we evaluate the performance per area metrics and present the results in Figure 17. As we can observe, FlexiBit provides better performance per area, 28% and 34% higher than Tensor Core and Bit-Fusion on average, respectively. The results indicate that benefits from flexible precision enabled by FlexiBit exceed hardware costs for supporting them.
@@ 5.3.4. Comparison against Bit-serial Accelerator
We select Cambricon-P(Hao et al., 2022) as the state-of-the-art bit-serial architecture, which also provides full flexibility in bit precisions. To consider both costs and benefits, we compare FlexiBit and Cambricon-P using Energy-Delay Product (EDP) in Figure 18. Our evaluation shows that the bit-serial architecture is area- and energy-efficient, which requires 1.8$\times$ less area and 3$\times$ less energy, on average, for GPT-3 on a Cloud-B accelerator. However, bit-serial architecture requires 12.6$\times$ more latency, which is a significant drawback. This led to 4.2$\times$ less EDP of FlexiBit, which demonstrates that the benefits exceed the costs.
@@ 5.3.5. Area Analysis
We present the area breakdown of FlexiBit and its PE in Figure 19 based on the Mobile-A configuration. We observe that core modules for flexible precision in FlexiBit’s PE, FBRT and Primitive Generator, account for about 50% of PE area. Although this implies the costs for supporting flexible precision, the performance benefits outweighs the costs as we discussed in Subsection 5.3.3 in the accelerator level.
@@ 6. Related Works
Table 5. Architecture categories and their readiness for LLM requirements with flexible precision/format FP/INT quantization.
@@
Flexible Precision Accelerators. Cambricon-P (Hao et al., 2022) processed different bit-serial bitflows in parallel, with bit-indexed Inner-Product Units to address the large dependency chains that cause the low hardware utilization. However, bit-serial architectures often suffer from limited performance when targeting a large-scale system. Also, Cambricon-P’s main target is scientific applications with a relatively small number of operands (compared to deep learning) with very long precision ($10^{3}$ $10^{5}$ bits), while deep learning typically uses billions of less than 16-bit data. Bit-fusion (Sharma et al., 2018) proposes fusible units for integer multiplication operation that enables utilizing all the logic for combinations of power-of-two operands (e.g. A8W4). Mix-GEMM (Reggiani et al., 2023) is another integer-only flexible-precision accelerator. RaPiD (Venkataramani et al., 2021) supports five different FP/INT precision/formats (FP16, FP8-e4m3, FP8-e5m2, FP9, INT4, and INT2). Although RaPiD introduces a novel architecture that maximizes usage of hardware components, the flexibility is limited to the five different presets. Following RaPiD, All-rounder (Noh et al., 2023b) supports more data precisions/formats. However, its performance degrades due to the costs for more flexibility. (Fu et al., 2021) and (Lebedev et al., 2014) support dynamic precision level adjustment with a trade-off between latency and accuracy. However, those works do not fully exploit precision adjustments for all types of AI models and tasks. As summarized in Table 5, only this work satisfies all the desired features for large language models with flexible FP/INT precision/format data.

### v2

@@ 5.3.1. Latency and Energy Comparison
Figure 10 presents the latency evaluation results of FlexiBit and baselines. We observe minor improvements for FP16-based models. However, as we deploy non-power-of-two precisions, FlexiBit significantly outperforms all the baselines, providing 59% and 31% less latency compared to Tensor Core and Bit-Fusion, on average. The improvements mainly originate from the capability to support arbitrary precision with high hardware underutilization. Baselines need to up-cast non-power-of-two precision data to the nearest power-of-two precision data, which leads to significant redundant bit operations. Another source of improvement is our $BitPacking$ technique which fully utilizes the SRAM capacity, which reduces DRAM traffic. To better demonstrate the impact of $BitPacking$, we study end-to-end latency of FlexiBit with and without $BitPacking$ scheme, as presented in Figure 11. $BitPacking$ improve the latency by 26%, on average, while the area and power overhead is negligible to the entire accelerator, as Subsection 5.3.4 presents.
@@
In terms of energy consumption, FlexiBit with weight-stationary dataflow provides considerably low energy, 66% and 33% lower than the Tensor Core and Bit-Fusion baselines, on average. The main source of energy saving is reduced on-chip traffic by utilizing packed memory layout. As a result, the energy saving is significant when non-power-of-two precision data is involved.
Figure 11. Latency comparison of FlexiBit with and without Bit-Packing structure. The results are normalized based on TensorCore Latency at each data precision.
@@
Figure 12. Performance per area comparison on 13 Act./Wgt. precision pairs. The sequence length is 2048.
@@
Figure 13. Comparison of a bit-serial (Cambricon-P and BitMod) and our bit-parallel (FlexiBit) flexible precision architecture using EDP normalized against Tensor Core-like baseline.
@@
Figure 14. The impact of register width and area breakdown. (a) PE Area breakdown of FlexiBit and throughput per area on the register width from 16 to 32 on Mobile-A scale (b) accelerator area breakdown (Reg_Width = 24)
@@ 5.3.2. Performance Per Area
To understand the relative benefits over costs, we evaluate the performance per area metrics and present the results in Figure 12. As we can observe, FlexiBit provides better performance per area, 28% and 34% higher than Tensor Core and Bit-Fusion on average, respectively. The results indicate that benefits from flexible precision enabled by FlexiBit exceed hardware costs for supporting them. However, TensorCore slightly outperforms FlexiBit in some power-of-two cases (i.e. [8, 8] and [4, 4]). We note that the main strength of FlexiBit is non-power-of-two precision computation, while it provides similar throughput for power-of-two precisions.
@@ 5.3.3. Comparison against Bit-serial Accelerator
We select Cambricon-P(Hao et al., 2022) and BitMod (Chen et al., 2025) as the state-of-the-art bit-serial architectures, which provide flexibility in bit precisions to some extent. We also utilize BitMod (Chen et al., 2025) simulator fo simulating BitMod with various precisions on weights. To consider both costs and benefits, we compare FlexiBit, Cambricon-P, and BitMod using Energy-Delay Product (EDP) in Figure 13. Our evaluation shows that Cambricon-P is best in power and energy-efficiency, working with 7.1$\times$ less power, on average. However, bit-serial architectures cannot achieve bit-parallel performance. Cambricon-P requires 52$\times$ more latency on Llama2-70b and Cloud-B scale due to the serialized computation for different precisions, which is a significant drawback. This led to 2.48$\times$ more EDP over Cambricon-P compared to FlexiBit on Llama-2-70b, demonstrating that the benefits of FlexiBit exceed the costs. Compared to BitMod, our accelerator provided 7.9$\times$ better latency. The significant latency differences originate from BitMod’s fixed precision for activations, long latencies for multiplications with larger bit widths, and the limited degree of bit parallelism. Table 4 and  Table 5 includes the performance and architecture details of all three accelerators. However, BitMod provides 2.7$\times$ higher energy efficiency due to its light-weight bit-serial mechanism. Latency and energy together, FlexiBit provides 2.9$\times$ lower EDP compared to BitMod to run Llama-2-70b model, which demonstrates that FlexiBit is a promising solution. In particular, the high performance is crucial for recent large-scale models, which makes FlexiBit a future-proof architecture.
Table 4. Average latency, energy, and EDP of different scales and accelerators on Llama-2-7b and Llama-2-70b models.
Scale Accel. Latency ($s$) Energy ($\mu J$) EDP ($\mu J.s$)
Llama2 Llama2 Llama2 Llama2 Llama2 Llama2
7b 70b 7b 70b 7b 70b
Mobile-B Cambricon-P(Hao et al., 2022) 83.95 1195.55 0.54 7.62 45.33 9110.09
BitMod(Chen et al., 2025) 6.94 151.12 2.99 36.18 20.75 5467.52
FlexiBit 1.52 20.52 9.84 135.86 14.95 2787.84
Cloud-B Cambricon-P(Hao et al., 2022) 20.58 249.28 0.38 4.65 7.82 1159.15
BitMod(Chen et al., 2025) 1.73 37.78 2.99 36.18 5.17 1366.88
FlexiBit 0.45 4.78 8.92 97.70 4.01 467.01
Table 5. Area and Power comparison with Mobile-A scale.
Scale Accel. Area ($mm^{2}$) Power (mW)
Mobile-A Cambricon-P(Hao et al., 2022) 5.11 122.15
Mobile-A BitMod(Chen et al., 2025) 4.70 629.76
Mobile-A FlexiBit 18.62 873.48
@@ 5.3.4. Area Analysis
We present the area breakdown of FlexiBit and its PE in Figure 14 based on the Mobile-A configuration. In the accelerator-level, we observe the negligible overhead of Bit-packing unit. Bit-packing implementation includes a base with 64-to-64 crossbar accompanied by an indexing mechanism for a 64-bit off-chip channel. For larger channel widths, area overhead doesn’t exponentially increase since we duplicate the base implementation (e.g., two base bit-packing units for 128-bit channel). We measure the same routing overhead as TensorCore in accelerator-level(12%), since FlexiBit is very similar to typical ML accelerators in high-level. In the PE-level, however, there is 6% routing and wiring overhead. We observe that core modules for flexible precision, FBRT and Primitive Generator, account for about 50% of PE area.
Determining the design parameter. We conduct an area analysis of FlexiBit’s PE varying the key design parameter, $reg\_width$. As the results in Figure 14(a) show, larger $reg\_width$ super-linearly increase. However, the relative overhead of crossbars does not meaningfully change. To determine proper $reg\_width$, we consider both area and performance utilizing throughput per area. As shown in Figure 14(a), we observe the best result for $reg_{w}idth=24$, which drove our design choice.
@@ 6. Related Works
Table 6. Architecture categories and their readiness for LLM requirements with flexible bit FP/INT quantization.
@@
Flexible Precision Accelerators. Cambricon-P (Hao et al., 2022) processed different bit-serial bitflows in parallel, with bit-indexed Inner-Product Units to address the large dependency chains that cause the low hardware utilization. BitMoD (Chen et al., 2025) introduces new precision representations that need dequantization on the fly. It uses multiple bit-serial multiplication lanes to perform W4A16 matrix multiplication and dequantization. However, bit-serial architectures often suffer from limited performance when targeting a large-scale system. Also, Cambricon-P’s main target is scientific applications with a relatively small number of operands (compared to deep learning) with very long precision ($10^{3}$ $10^{5}$ bits), while deep learning typically uses billions of less than 16-bit data. Bit-fusion (Sharma et al., 2018) proposes fusible units for integer multiplication operation that enables utilizing all the logic for combinations of power-of-two operands (e.g. A8W4). Mix-GEMM (Reggiani et al., 2023) is another integer-only flexible-precision accelerator. RaPiD (Venkataramani et al., 2021) supports five different FP/INT precision/formats (FP16, FP8-e4m3, FP8-e5m2, FP9, INT4, and INT2). Although RaPiD introduces a novel architecture that maximizes usage of hardware components, the flexibility is limited to the five different presets. Following RaPiD, All-rounder (Noh et al., 2023b) supports more data precisions/formats. However, its performance degrades due to the costs for more flexibility. (Fu et al., 2021) and (Lebedev et al., 2014) support dynamic precision level adjustment with a trade-off between latency and accuracy. However, those works do not fully exploit precision adjustments for all types of AI models and tasks. As summarized in Table 6, only this work satisfies all the desired features for large language models with flexible FP/INT precision/format data.